Mr. paddle took his boat 80 miles down a river and back.The trip downsteam took 4hrs while back took 8hrs.find the speed of the boat in still water and the speed of the current?

Answer by ptfile(81)

Let 80 miles = the distance Mr. Paddle traveled
4hrs = the duration of the boat traveled on downstream
8hrs = the duration of the boat traveled on upstream
Mr. paddle took his boat 80 miles down a river and back. The trip downstream took 4hrs while back took 8hrs.
Distance = (rate)(time)
rate = Distance/time
80 = (x+y)4
10 = x+y
80 = (x-y)8
20 = x-y
x+y = 20
x-y = 10
2x = 30
x = 15
Substitute x with 15
15+y = 20
y = 5
15-y = 10
y = 5
The boat's speed is 15kph while the current's speed is 5kph.